OpenStudy (anonymous):

Help: What is the quotient of 4x^3 + 5x^2 - 12 + 27/(x + 3)? A. 4x^2 - 7x + 9 B. 4x^2 + 7x + 9 C. 4x^2 - 7x + 27 D. 4x^2 + 7x + 27 I think it's C.

OpenStudy (anonymous):

umm hold on

OpenStudy (anonymous):

@bohotness she is really smart

OpenStudy (bohotness):

Final result : (3x2 - 2x + 1) • (x + 2) • -1 Step by step solution : Step 1 : Raise x to the 2nd power Exponentiaion : Equation at the end of step 1 : ((((((6•(x6))+(20•(x5)))+(7•(x4)))+(27•(x2)))-14x)+8) ————————————————————————————————————————————————————— ((((0-(2•(x3)))-(4•x2))+x)-4) Step 2 : Raise x to the 3rd power Exponentiaion : Equation at the end of step 2 : ((((((6•(x6))+(20•(x5)))+(7•(x4)))+(27•(x2)))-14x)+8) ————————————————————————————————————————————————————— ((((0-(2•x3))-22x2)+x)-4) Step 3 : Raise x to the 2nd power Exponentiaion : Equation at the end of step 3 : ((((((6•(x6))+(20•(x5)))+(7•(x4)))+(27•x2))-14x)+8) ——————————————————————————————————————————————————— (-2x3-4x2+x-4) Step 4 : Raise x to the 4th power Exponentiaion : Equation at the end of step 4 : ((((((6•(x6))+(20•(x5)))+(7•x4))+33x2)-14x)+8) —————————————————————————————————————————————— (-2x3-4x2+x-4) Step 5 : Raise x to the 5th power Exponentiaion : Equation at the end of step 5 : ((((((6•(x6))+(20•x5))+7x4)+33x2)-14x)+8) ————————————————————————————————————————— (-2x3-4x2+x-4) Step 6 : Raise x to the 6th power Exponentiaion : Equation at the end of step 6 : ((((((6•x6)+(22•5x5))+7x4)+33x2)-14x)+8) ———————————————————————————————————————— (-2x3-4x2+x-4) Step 7 : 6x6 + 20x5 + 7x4 + 27x2 - 14x + 8 Simplify ————————————————————————————————— -2x3 - 4x2 + x - 4 Trying to factor by pulling out : 7.1 Factoring: 6x6 + 20x5 + 7x4 + 27x2 - 14x + 8 Thoughtfully split the expression at hand into groups, each group having two terms : Group 1: 7x4 + 27x2 Group 2: 6x6 + 20x5 Group 3: -14x + 8 Pull out from each group separately : Group 1: (7x2 + 27) • (x2) Group 2: (3x + 10) • (2x5) Group 3: (7x - 4) • (-2) Looking for common sub-expressions : Group 1: (7x2 + 27) • (x2) Group 3: (7x - 4) • (-2) Group 2: (3x + 10) • (2x5) Bad news !! Factoring by pulling out fails : The groups have no common factor and can not be added up to form a multiplication. Pulling out like terms : 7.2 Pull out like factors : -2x3 - 4x2 + x - 4 = -1 • (2x3 + 4x2 - x + 4) Checking for a perfect cube : 7.3 2x3 + 4x2 - x + 4 is not a perfect cube Trying to factor by pulling out : 7.4 Factoring: 2x3 + 4x2 - x + 4 Thoughtfully split the expression at hand into groups, each group having two terms : Group 1: -x + 4 Group 2: 4x2 + 2x3 Pull out from each group separately : Group 1: (-x + 4) • (1) = (x - 4) • (-1) Group 2: (x + 2) • (2x2) Bad news !! Factoring by pulling out fails : The groups have no common factor and can not be added up to form a multiplication. Polynomial Roots Calculator : 7.5 Find roots (zeroes) of : F(x) = 6x6 + 20x5 + 7x4 + 27x2 - 14x + 8 Polynomial Roots Calculator is a set of methods aimed at finding values of x for which F(x)=0 Rational Roots Test is one of the above mentioned tools. It would only find Rational Roots that is numbers x which can be expressed as the quotient of two integers The Rational Root Theorem states that if a polynomial zeroes for a rational number P/Q then P is a factor of the Trailing Constant and Q is a factor of the Leading Coefficient In this case, the Leading Coefficient is 6 and the Trailing Constant is 8. The factor(s) are: of the Leading Coefficient : 1,2 ,3 ,6 of the Trailing Constant : 1 ,2 ,4 ,8 Let us test .... P Q P/Q F(P/Q) Divisor -1 1 -1.00 42.00 -1 2 -0.50 21.66 -1 3 -0.33 15.68 -1 6 -0.17 11.09 -2 1 -2.00 0.00 x + 2 Note - For tidiness, printing of 15 checks which found no root was suppressed The Factor Theorem states that if P/Q is root of a polynomial then this polynomial can be divided by q*x-p Note that q and p originate from P/Q reduced to its lowest terms In our case this means that 6x6 + 20x5 + 7x4 + 27x2 - 14x + 8 can be divided with x + 2 Polynomial Long Division : 7.6 Polynomial Long Division Dividing : 6x6 + 20x5 + 7x4 + 27x2 - 14x + 8 ("Dividend") By : x + 2 ("Divisor") dividend 6x6 + 20x5 + 7x4 + 27x2 - 14x + 8 - divisor * 6x5 6x6 + 12x5 remainder 8x5 + 7x4 + 27x2 - 14x + 8 - divisor * 8x4 8x5 + 16x4 remainder - 9x4 + 27x2 - 14x + 8 - divisor * -9x3 - 9x4 - 18x3 remainder 18x3 + 27x2 - 14x + 8 - divisor * 18x2 18x3 + 36x2 remainder - 9x2 - 14x + 8 - divisor * -9x1 - 9x2 - 18x remainder 4x + 8 - divisor * 4x0 4x + 8 remainder 0 Quotient : 6x5+8x4-9x3+18x2-9x+4 Remainder: 0 Polynomial Roots Calculator : 7.7 Find roots (zeroes) of : F(x) = 6x5+8x4-9x3+18x2-9x+4 See theory in step 7.5 In this case, the Leading Coefficient is 6 and the Trailing Constant is 4. The factor(s) are: of the Leading Coefficient : 1,2 ,3 ,6 of the Trailing Constant : 1 ,2 ,4 Let us test .... P Q P/Q F(P/Q) Divisor -1 1 -1.00 42.00 -1 2 -0.50 14.44 -1 3 -0.33 9.41 -1 6 -0.17 6.05 -2 1 -2.00 102.00 -2 3 -0.67 21.46 -4 1 -4.00 -3192.00 -4 3 -1.33 69.33 1 1 1.00 18.00 1 2 0.50 3.56 1 3 0.33 2.79 1 6 0.17 2.97 2 1 2.00 306.00 2 3 0.67 5.70 4 1 4.00 7872.00 4 3 1.33 53.23 Polynomial Roots Calculator found no rational roots Polynomial Roots Calculator : 7.8 Find roots (zeroes) of : F(x) = 2x3+4x2-x+4 See theory in step 7.5 In this case, the Leading Coefficient is 2 and the Trailing Constant is 4. The factor(s) are: of the Leading Coefficient : 1,2 of the Trailing Constant : 1 ,2 ,4 Let us test .... P Q P/Q F(P/Q) Divisor -1 1 -1.00 7.00 -1 2 -0.50 5.25 -2 1 -2.00 6.00 -4 1 -4.00 -56.00 1 1 1.00 9.00 1 2 0.50 4.75 2 1 2.00 34.00 4 1 4.00 192.00 Polynomial Roots Calculator found no rational roots Polynomial Long Division : 7.9 Polynomial Long Division Dividing : 6x5+8x4-9x3+18x2-9x+4 ("Dividend") By : 2x3+4x2-x+4 ("Divisor") dividend 6x5 + 8x4 - 9x3 + 18x2 - 9x + 4 - divisor * 3x2 6x5 + 12x4 - 3x3 + 12x2 remainder - 4x4 - 6x3 + 6x2 - 9x + 4 - divisor * -2x1 - 4x4 - 8x3 + 2x2 - 8x remainder 2x3 + 4x2 - x + 4 - divisor * x0 2x3 + 4x2 - x + 4 remainder 0 Quotient : 3x2-2x+1 Remainder : 0 Trying to factor by splitting the middle term 7.10 Factoring 3x2-2x+1 The first term is, 3x2 its coefficient is 3 . The middle term is, -2x its coefficient is -2 . The last term, "the constant", is +1 Step-1 : Multiply the coefficient of the first term by the constant 3 • 1 = 3 Step-2 : Find two factors of 3 whose sum equals the coefficient of the middle term, which is -2 . -3 + -1 = -4 -1 + -3 = -4 1 + 3 = 4 3 + 1 = 4 Observation : No two such factors can be found !! Conclusion : Trinomial can not be factored Final result : (3x2 - 2x + 1) • (x + 2) • -1